# Model-Free Assessment of Simulator Fidelity via Quantile Curves

### v2

@@ Abstract
As generative AI models are increasingly used to simulate real-world systems, quantifying the “sim-to-real” gap is critical. The distributional discrepancy between real and simulated outputs is a random variable driven by the stochastic input scenario. A fundamental challenge is that for any given input, the ground-truth and simulated output distributions are only observable through finite batches of samples, often of heterogeneous sizes. This renders standard predictive inference methods inapplicable, as they seek to quantify uncertainty in observable outputs rather than their underlying population parameters. To address this, we construct confidence sets for these latent parameters and use them to derive a robust proxy for the sim-to-real discrepancy. We then estimate the quantile function of this proxy to provide a comprehensive risk profile of the simulator. Our method is model-agnostic and handles general output spaces, such as categorical survey responses and continuous multi-dimensional sensor data. By rigorously accounting for sampling error, the resulting risk profile supports statistical inference for the real output distribution in a new scenario, the calculation of risk measures like Conditional Value-at-Risk (CVaR), and principled comparisons across simulators. We demonstrate the practical utility of this method by evaluating the alignment of four major LLMs with human populations on the WorldValueBench dataset.
@@ 1 Introduction
Against this backdrop, quantifying the distributional discrepancy between simulated and real-world outcomes is critical. This “sim-to-real” gap is a pervasive challenge in domains ranging from LLMs to computing systems and robotics (Gao et al., 2025; Durmus et al., 2023; Tobin et al., 2017; Peng et al., 2018). Crucially, this gap is not static; it is a function of the scenario (e.g., operating conditions or user prompts). Because a simulator is deployed across a heterogeneous population of scenarios, the sim-to-real gap behaves as a random variable induced by the underlying scenario distribution. Consequently, assessing fidelity requires characterizing the distribution of this random variable, rather than merely reporting a single average error. This is the primary objective of our work.
@@
In contrast, modern digital twins rely on complex ML models (e.g., an LLM or a deep neural network), and one often does not have access to the model directly. Moreover, it can be computationally prohibitive to calibrate its parameters exactly. Consequently, we treat the simulator as a black box and focus on directly characterizing the sim-to-real discrepancy of the output—an approach termed output uncertainty quantification by Jeon et al. (2024). A further distinction is that much of the UQ literature is asymptotic, whereas we provide finite-sample guarantees, which is critical when only a limited number of scenarios are available. Finally, while the classical input uncertainty literature is primarily concerned with Monte Carlo noise, which is the variance induced by stochastic sampling conditional on a fixed input model, our work incorporates both variance and bias of the simulator relative to the real world. This is precisely the focus of “sim-to-real” discrepancy.
Recent work in the LLM literature has also begun to study output uncertainty using model-agnostic discrepancy measures. For example, Santurkar et al. (2023a) aggregate the sim-to-real gap into a single scalar summary of overall error or bias, and Huang et al. (2025) bound the discrepancy at a fixed quantile level. These approaches correspond to evaluating particular functionals of the distribution of the sim-to-real discrepancy. More generally, the literature has largely emphasized scalar summaries or fixed-level bounds, leaving open the question of how to obtain a calibrated, distribution-level risk profile across scenarios.
@@ 2 Framework and Motivating Example
We begin with a motivating use case and then formalize the quantile estimation problem; Appendix D gives additional examples.
@@
Let $p(\psi)$ be a population statistic of interest, which is a functional of the conditional distribution $Q^{\mathrm{gt}}(\cdot\mid\psi)$ and lives in a parameter space $\Theta$. In our running example, $\Theta=\mathcal{P}^{d}$, $p(\psi)$ is the mean response of survey respondents, and simulator can be defined similarly under the simulator population $\mathcal{P}^{\mathrm{sim}}$. Further examples of $p(\psi)$ are given in Section 3. To simplify notation, we will write $p(\psi)$ and $q(\psi)$ as $p_{\psi}$ and $q_{\psi}$. Summing up:
@@
Construct a function $\hat{V}(\cdot,\mathcal{D}):(0,1)\to\mathbb{R}$ such that, for a new $\psi\sim\Psi$ and all $\alpha\in(0,1)$,
$\mathbb{P}_{\psi\sim\Psi}\big(\Delta_{\psi}\leq\hat{V}(\alpha,\mathcal{D})\big|\mathcal{D}\big)\geq\alpha-\varepsilon_{m},$
@@
Note that $\hat{V}$ is indexed by both the prescribed level $\alpha$ and the dataset $\mathcal{D}$. Establishing our desired result requires controlling two sources of uncertainty: (i) scenario uncertainty, since we observe only finitely many scenarios $\{\psi_{j}\}_{j=1}^{m}\sim\Psi$, and (ii) finite-sample uncertainty within each scenario, since $p_{j}$ and $q_{j}$ are only observed through finite samples that produce $\hat{p}_{j}$ and $\hat{q}_{j}$. Our construction of $\hat{V}(\alpha,\mathcal{D})$ explicitly accounts for both, and the main theorem will yield non-vacuous, finite-sample coverage guarantees.
@@ 3 Theoretical Result
In this section, we construct $\hat{V}$, establish its theoretical guarantees, and conclude by discussing the benefits of distribution-level fidelity guarantees.
@@ 3.1 Methodology
    $\hat{V}_{m}\!\left(\alpha\right):=\text{the $\alpha$-quantile of}\{\hat{\Delta}_{j}\}_{j=1}^{m}.$ (3.3)
We will show that $\hat{V}_{m}\!\left(\alpha\right)$ is a calibrated quantile curve for the discrepancy distribution, up to explicit finite-sample corrections. Note that the key design choice in our method is the confidence set $\mathcal{C}_{j}\!\left(\cdot\right)$ that quantifies the uncertainty of $\hat{p}_{j}$. This uncertainty set is the key device that links per-scenario sampling error to distribution-level calibration. Next, we show how to construct these sets using off-the-shelf concentration bounds, starting with the multinomial setting.
@@ 3.2 Calibrated Quantile Curve Theory
Before stating our main result, we consider the special case $\gamma_{j}\equiv\tfrac{1}{2}$. Then, in the asymptotic regime $m\to\infty$, our guarantee (3.6) below roughly implies the following: for any target quantile level $\tau\in(0,1)$,
@@
In general, the $\gamma_{j}$’s can be different, and we obtain the following result:
@@
Suppose Assumptions 3.13.2 hold and $\gamma_{j}\in(0,1)$ for each $j\in[m]$. Let $\mathcal{C}_{j}(\hat{p}_{j},\gamma_{j})$ denote any $\gamma_{j}$-confidence set (see (3.1)), $\hat{\Delta}_{j}$ denote the pseudo-discrepancy (see (3.2)), and let $\hat{V}_{m}$ denote the quantile function of the pseudo-discrepancy (see (3.3)). For $\alpha\in(0,1)$ let $\alpha]_{m}:=m^{-1}\lceil m\alpha\rceil\in\Big[\alpha,\alpha+\frac{1}{m}\Big]$ and set
@@ 3.3 Benefits of Distribution-Level Characterization
In addition, the calibrated quantile curve can be compressed to provide summary statistics. Returning to the motivating example, a media research firm can use a calibrated AUC to assess average simulator bias, while a pollster or risk-averse product team may track a calibrated $\mathrm{CVaR}_{\alpha}$ to bound rare but consequential misreads of public sentiment.
@@ 4.1 Dataset and Methodology
Since the outcome is bounded in [1,1][-1,1], we construct per-scenario confidence sets $\mathcal{C}_{j}$ using Example 3.2. We set $\gamma_{j}=1-n_{j}^{-1/3},$ which is increasing in sample size $n_{j}$ [^2] As discussed in Section 3, the coverage level $\gamma_{j}=g(n_{j})$ should increase with the sample size $n_{j}$. We therefore adopt a generic form $\gamma_{j}=1-n_{j}^{-\beta},$ and conduct a sensitivity analysis on $\beta$ value in Appendix C.4., and yields an average coverage level of $\bar{\gamma}_{m}\approx 0.87$. The loss function $L(p,q)=(p-q)^{2}$. Next, we use the procedure described in Section 3 (see (3.1)-(3.3)) to obtain a fidelity profile for each candidate LLM, and plot the calibrated curve using the asymptotic index adjustment in (3.5). For completeness, we also apply the same workflow to a Bernoulli setting (EEDI) and a multinomial setting (OpinionQA); see Appendix E.
@@ 4.2 Fidelity Profiling
Figure 3: Calibrated $\hat{V}(\tau)$ across LLMs.
The calibrated quantile functions for the four different LLMs are plotted in Figure 3. Lower, flatter curves indicate uniformly small discrepancies, while elbows reveal rare but severe misses. Our results suggest that all LLM-simulators outperform the uniform benchmark on over $80\%$ of questions, but fail to capture the tail scenarios. GPT-4o is the lowest across all quantiles, indicating the most reliable alignment, with GPT-5-mini close behind yet failing to capture some outlier questions, and Llama 3.3 and QWen-3 trail behind.
Figure 4: Robustness check of simulator performances.
@@ 5 Tightness Analysis of Calibrated Quantiles
Figure 5: Tightness analysis for varying $n_{j}$ under GPT-4o.
@@
The results in Figure 5 correspond to $\gamma_{j}=1-n_{j}^{-\frac{1}{3}}$. In order to isolate the impact of batch-size-specific $\gamma_{j}$, as we discussed in Section 3, we repeat the same experiment by setting $\gamma_{j}\equiv\tfrac{1}{2}$ for all questions. The corresponding results are plotted in Figure 6. The plots here exhibit a persistent vertical separation between the oracle quantile curve and the calibrated curves, even for large $n_{j}$. This phenomenon is explained by the fact that here $\bar{\gamma}_{m}=\tfrac{1}{2}$ (independent of $n_{j}$) and the asymptotic calibration evaluates the pseudo-quantile at $\hat{V}_{m}\big(0.5\tau+0.5\big)$. Consequently, at lower quantile levels ($\tau\approx 0$), the our procedure approximates the quantile function by at least the median, inducing a non-vanishing conservativeness that cannot be eliminated by increasing $n_{j}$. This underscores the importance of allowing scenario-dependent coverage levels.
Figure 6: Tightness analysis for varying $n_{j}$ under GPT-4o with fixed $\gamma_{j}=\tfrac{1}{2}$.
@@
Figure 7: Confidence Bands of GPT-4o and Llama.

### v3

@@ Abstract
As generative AI models are increasingly used to simulate real-world systems, quantifying the “sim-to-real” gap is critical. For each input setting of interest—which we call a scenario, such as a survey question or operating condition—the real and simulated systems are associated with unobserved latent population parameters, and their discrepancy varies across scenarios. A fundamental challenge is that, for any given scenario, this discrepancy cannot be observed directly, since both systems are accessible only through finite samples, often of heterogeneous sizes across scenarios. Standard predictive inference methods are therefore ill-suited, as they quantify uncertainty in observable outputs rather than latent population parameters. To address this, we construct confidence sets for these latent parameters and use them to derive a robust proxy for the sim-to-real discrepancy. We then estimate the quantile function of this proxy to obtain a distribution-level risk profile of the simulator, which supports a broad range of statistical summaries, including statistical inference for the real output distribution in a new scenario, the calculation of risk measures like Conditional Value-at-Risk (CVaR), and principled comparisons across simulators. Our method is model-agnostic and handles general output spaces, such as categorical survey responses and continuous multi-dimensional data. We demonstrate the practical utility of this method by evaluating the alignment of four major LLMs with human populations on the WorldValueBench dataset.
@@ 1 Introduction
Against this backdrop, quantifying the distributional discrepancy between simulated and real-world outcomes is critical. This “sim-to-real” gap is a pervasive challenge in domains ranging from LLMs to computing systems and robotics (Gao et al., 2025; Durmus et al., 2023; Tobin et al., 2017; Peng et al., 2018). Crucially, it depends on the scenario (e.g., operating conditions or user prompts). In our setting, each scenario is associated with an unobserved latent population quantity for both the real system and the simulator, observed only through finite samples. As a simulator is deployed over a heterogeneous population of scenarios, the discrepancy between these latent scenario-level quantities becomes a random variable induced by the scenario distribution. Assessing fidelity therefore requires characterizing this discrepancy distribution, rather than a single average error, which can hide rare but consequential failures. This distributional characterization is the primary objective of our work.
@@
In contrast, modern digital twins rely on complex ML models (e.g., an LLM or a deep neural network), and one often does not have access to the model directly. Moreover, it can be computationally prohibitive to calibrate its parameters exactly. Consequently, we treat the simulator as a black box and focus on directly characterizing the sim-to-real discrepancy of the output—an approach termed output uncertainty quantification by Jeon et al. (2024). This perspective is related to simulation-based inference (SBI) (Cranmer et al., 2020; Papamakarios et al., 2019), which likewise studies inference with complex simulators, often under implicit generative models. The key difference is that SBI typically targets posterior inference on latent simulator parameters, whereas our goal is to assess the output-level fidelity of a fixed simulator across scenarios. Returning to the broader UQ literature, another distinction is that much of the work is asymptotic, whereas we provide finite-sample guarantees, which is critical when only a limited number of scenarios are available. Finally, while the classical input uncertainty literature is primarily concerned with Monte Carlo noise, which is the variance induced by stochastic sampling conditional on a fixed input model, our work incorporates both variance and bias of the simulator relative to the real world. This is precisely the focus of “sim-to-real” discrepancy.
Recent work in the LLM literature has also begun to study output uncertainty using model-agnostic discrepancy measures. For example, Santurkar et al. (2023a) aggregate the sim-to-real gap into a single scalar summary of overall error or bias, and Huang et al. (2025) bound the discrepancy at a fixed quantile level. These approaches correspond to evaluating particular functionals of the distribution of the sim-to-real discrepancy. More generally, the literature has largely emphasized scalar summaries or fixed-level bounds, leaving open the question of how to obtain a calibrated, distribution-level risk profile across scenarios. Two-sample testing is another nearby line of work. Kernel and related two-sample tests (Gretton et al., 2012; Schrab et al., 2023) are similarly sample-based and often model-agnostic, but their primary goal is to determine whether two samples are distinguishable or consistent with having been drawn from the same distribution. By contrast, our target is not a global equality test between pooled real and simulated data, but the calibrated quantile profile of scenario-level sim-to-real discrepancy across scenarios.
@@ 2 Framework and Motivating Example
We begin with a motivating use case through a survey-style LLM and then formalize the quantile estimation problem. Although we introduce the notation through an LLM example, the framework itself is more general. In Appendix D, we illustrate how the same formulation applies to other settings, including discrete-event manufacturing and broader control/scientific simulators with categorical and continuous outputs. Specifically, we ran our main numerical analysis under a survey setting, but conducted additional analysis in Appendix E under Building control setting in Appendix D.
@@
Let $p(\psi)$ be a population statistic of interest, which is a functional of the conditional distribution $Q^{\mathrm{gt}}(\cdot\mid\psi)$ and lives in a parameter space $\Theta$. In our running example, $\Theta=\mathcal{P}^{d}$, $p(\psi)$ is the mean response of survey respondents, and simulator statistic of interest $q(\psi)$ can be defined similarly under the simulator population $\mathcal{P}^{\mathrm{sim}}$. Further examples of $p(\psi)$ are given in Section 3. To simplify notation, we will write $p(\psi)$ and $q(\psi)$ as $p_{\psi}$ and $q_{\psi}$. Summing up:
@@
Construct a function $\hat{V}_{m}(\cdot,\mathcal{D}):(0,1)\to\mathbb{R}$ such that, for a new $\psi\sim\Psi$ and all $\tau\in(0,1)$,
$\mathbb{P}_{\psi\sim\Psi}\big(\Delta_{\psi}\leq\hat{V}_{m}(\tau,\mathcal{D})\big|\mathcal{D}\big)\geq\tau-\varepsilon_{m},$
@@
Note that $\hat{V}_{m}$ is indexed by both the prescribed level $\tau$ and the dataset $\mathcal{D}$. Establishing our desired result requires controlling two sources of uncertainty: (i) scenario uncertainty, since we observe only finitely many scenarios $\{\psi_{j}\}_{j=1}^{m}\sim\Psi$, and (ii) finite-sample uncertainty within each scenario, since $p_{j}$ and $q_{j}$ are only observed through finite samples that produce $\hat{p}_{j}$ and $\hat{q}_{j}$. Our construction of $\hat{V}_{m}(\tau,\mathcal{D})$ explicitly accounts for both, and the main theorem will yield non-vacuous, finite-sample coverage guarantees.
We summarize the notations introduced in this section in Table 2, along with few upcoming notations from Section 3.
Table 2: Notation summary.
Symbol Meaning
$\psi\sim\Psi$ Scenario drawn from the scenario population.
$z\sim P$ Real-world respondent drawn from the population distribution.
$p(\psi)$(or $p_{\psi}$) Real-world population parameter for scenario $\psi$.
$q(\psi)$(or $q_{\psi}$) Simulator-side population parameter for scenario $\psi$ under fixed simulator settings.
$\hat{p}_{\psi},\hat{q}_{\psi}$ Finite-sample estimators of $p_{\psi}$ and $q_{\psi}$.
$n_{j}$ Human sample size for scenario $j$.
$k$ (or $k_{j}$) Simulator sample size for scenario $j$.
$L(\cdot,\cdot)$ Discrepancy function between real and simulated distribution.
$\Delta_{\psi}$ True sim-to-real discrepancy for scenario $\psi$.
$\gamma_{j}$ Per-scenario confidence-set coverage level.
$\bar{\gamma}_{m}$ Average coverage level across $m$ scenarios.
$\mathcal{C}_{j}(\hat{p}_{j},\gamma_{j})$ $\gamma_{j}$-confidence set for the latent parameter $p_{j}$.
$\hat{\Delta}_{j}$ Pseudo-discrepancy for scenario $j$, defined via a confidence set around $\hat{p}_{j}$.
$\hat{V}_{m}(\tau)$ Pseudo-discrepancy quantile curve built from $m$ scenarios, indexed by quantile level $\tau$.
$\tau$ Quantile-function argument, counted in the usual direction from lower to upper quantiles.
$\alpha$ Upper-tail fraction / exceedance level used when discussing tail scenarios.
@@ 3 Theoretical Result
In this section, we construct $\hat{V}_{m},$, establish its theoretical guarantees, and conclude by discussing the benefits of distribution-level fidelity guarantees.
@@ 3.1 Methodology
    $\hat{V}_{m}\!\left(\tau\right):=\text{the $\tau$-quantile of}\{\hat{\Delta}_{j}\}_{j=1}^{m}.$ (3.3)
We will show that $\hat{V}_{m}\!\left(\tau\right)$ is a calibrated quantile curve for the discrepanwILLY624 cy distribution, up to explicit finite-sample corrections. Note that the key design choice in our method is the confidence set $\mathcal{C}_{j}\!\left(\cdot\right)$ that quantifies the uncertainty of $\hat{p}_{j}$. This uncertainty set is the key device that links per-scenario sampling error to distribution-level calibration. Next, we show how to construct these sets using off-the-shelf concentration bounds, starting with the multinomial setting.
@@ 3.2 Calibrated Quantile Curve Theory
Before stating our main result, we consider the special case $\gamma_{j}\equiv\tfrac{1}{2}$. Then, in the asymptotic regime $m\to\infty$, our guarantee (3.6) implies the following: for any target quantile level $\tau\in(0,1)$,
@@
In general, the $\gamma_{j}$’s can be different, and we obtain the following asymptotic result:
@@
Suppose Assumptions 3.13.2 hold and $\gamma_{j}\in(0,1)$ for each $j\in[m]$. Let $\mathcal{C}_{j}(\hat{p}_{j},\gamma_{j})$ denote any $\gamma_{j}$-confidence set (see (3.1)), $\hat{\Delta}_{j}$ denote the pseudo-discrepancy (see (3.2)), and let $\hat{V}_{m}$ denote the quantile function of the pseudo-discrepancy (see (3.3)). For $\alpha\in(0,1)$, let $\alpha]_{m}:=m^{-1}\lceil m\alpha\rceil$ denote the rounded empirical grid point associated with $\alpha$, so that $\alpha]_{m}\in[\alpha,\alpha+\frac{1}{m}]$, and let $x]_{+}:=\max\{x,0\}$. Set
@@ 3.3 Benefits of Distribution-Level Characterization
In addition, the calibrated quantile curve can be compressed to provide summary statistics, hence strictly more informative than any single risk summary. Returning to the motivating example, a media research firm can use a calibrated AUC to assess average simulator bias, while a pollster or risk-averse product team may track a calibrated $\mathrm{CVaR}_{\alpha}$ to bound rare but consequential misreads of public sentiment.
@@ 4.1 Dataset and Methodology
Since the outcome is bounded in [1,1][-1,1], we construct per-scenario confidence sets $\mathcal{C}_{j}$ using Example 3.2. We set $\gamma_{j}=1-n_{j}^{-1/3},$ which is increasing in sample size $n_{j}$ [^2] As discussed in Section 3, the coverage level $\gamma_{j}=g(n_{j})$ should increase with the sample size $n_{j}$. We therefore adopt a generic form $\gamma_{j}=1-n_{j}^{-\beta},$ and conduct a sensitivity analysis on $\beta$ value in Appendix C.4., and yields an average coverage level of $\bar{\gamma}_{m}\approx 0.87$. The loss function $L(p,q)=(p-q)^{2}$. Next, we use the procedure described in Section 3 (see (3.1)-(3.3)) to obtain a fidelity profile for each candidate LLM, and plot the calibrated curve using the asymptotic index adjustment in (3.5). For completeness, we also apply the same workflow to a Bernoulli setting (EEDI), multinomial setting (OpinionQA) and a continuous temperature prediction setting (UMAR dataset from NEST); see Appendix E.
@@ 4.2 Fidelity Profiling
Figure 3: Calibrated discrepancy quantile curves $\hat{V}(\tau)$ across LLM simulators. Colors: blue = GPT-4o, green = GPT-5-MINI, orange = LLAMA 3.3, purple = QWEN-3, and red = the uniform baseline.
The calibrated quantile functions for the four different LLMs are plotted in Figure 3. Lower, flatter curves indicate uniformly small discrepancies, while elbows reveal rare but severe misses. Our results suggest that all LLM-simulators outperform the uniform benchmark on over $80\%$ of questions, but fail to capture the tail scenarios. GPT-4o is the lowest across all quantiles, indicating the most reliable alignment, with GPT-5-mini close behind yet failing to capture some outlier questions, and Llama 3.3 and QWen-3 trail behind. While this comparison is useful, we want to emphasize that the uniform baseline is not intended to represent a universally poor simulator. For some near-balanced questions, it can return low sim-to-real gap. Its role here is instead to serve as a simple structure-free reference for illustrating how the proposed framework compares simulator fidelity at the scenario level.
Figure 4: Robustness check of simulator performance under different human sample sizes. Panels correspond to $n_{j}=50$ (top left), $n_{j}=500$ (top right), $n_{j}=5000$ (bottom left), and $n_{j}=10000$ (bottom right), with $k=200$ fixed. Colors match Figure 3.
@@ 5 Tightness Analysis of Calibrated Quantiles
Figure 5: Tightness analysis for varying $n_{j}$ under GPT-4o. The dotted black curve is the oracle quantile curve $\Delta^{\star}$; the colored curves are calibrated curves computed from subsamples with $n_{j}=50$ (blue), $n_{j}=500$ (orange), $n_{j}=1000$ (green), and $n_{j}=5000$ (red). The corresponding adaptive average coverage levels shown in the legend are $\bar{\gamma}\approx 0.73,0.87,0.90,$ and $0.94$, respectively.
@@
The results in Figure 5 correspond to $\gamma_{j}=1-n_{j}^{-\frac{1}{3}}$. In order to isolate the impact of batch-size-specific $\gamma_{j}$, as we discussed in Section 3, we repeat the same experiment by setting $\gamma_{j}\equiv\tfrac{1}{2}$ for all questions. The corresponding results are plotted in Figure 6. The plots here exhibit a persistent vertical separation between the oracle quantile curve and the calibrated curves, even for large $n_{j}$. This phenomenon is explained by the fact that here $\bar{\gamma}_{m}=\tfrac{1}{2}$ (independent of $n_{j}$) and the asymptotic calibration evaluates the pseudo-quantile at $\hat{V}_{m}\big(0.5\tau+0.5\big)$. Consequently, at lower quantile levels ($\tau\approx 0$), our procedure approximates the quantile function by at least the median, inducing a non-vanishing conservativeness that cannot be eliminated by increasing $n_{j}$. This underscores the importance of allowing scenario-dependent coverage levels.
Figure 6: Tightness analysis for varying $n_{j}$ under GPT-4o with fixed $\gamma_{j}=\tfrac{1}{2}$. Colors match Figure 5.
@@
Figure 7: Confidence bands for GPT-4o and LLAMA $3.$ The top panel uses $n_{j}=1000$, and $\bar{\gamma}_{L}=\bar{\gamma}_{U}=0.900$; the bottom panel uses $n_{j}=5000$, and $\bar{\gamma}_{L}=\bar{\gamma}_{U}=0.942$, both with $k=200$. In each panel, blue curves denote the lower and upper bands for GPT-4o, and orange curves for LLAMA 3.3 respectively.